OpenStudy (anonymous):

Simplify 3sqrt 400x^5y^2

OpenStudy (anonymous):

\[3\sqrt{400x^5y^2}\]

OpenStudy (luigi0210):

Break them up into simpler terms: \[\sqrt{4*100*x^4*x*y^2}\]

OpenStudy (anonymous):

I got the answer but i want to know if you'll get \[180x^2y \sqrt{y}\]

OpenStudy (anonymous):

my bad

OpenStudy (luigi0210):

Hm, you might wanna try again :)

OpenStudy (anonymous):

well the factors of 400 is 5*5*2*2*2*2. So you'll be left with \[5*3*2^2\sqrt{nothing here} is that what you'\ll get?\]

OpenStudy (luigi0210):

There's something under the rad, it's just not a y

OpenStudy (anonymous):

the two?

OpenStudy (luigi0210):

 Make sense?